OpenStudy (anonymous):

What is the product of -32 and -24?

OpenStudy (anonymous):

a) multiply b) use a calculator c) google

OpenStudy (anonymous):

multiply 32x24...then make it negative, try it.

OpenStudy (anonymous):

-32 * -24,

OpenStudy (anonymous):

actually the answer is positive

OpenStudy (anonymous):

768

OpenStudy (anonymous):

Kris :D

OpenStudy (anonymous):

Yes negative times negative is positive.

OpenStudy (anonymous):

ohh sorry it is, negative and negative is positive

OpenStudy (anonymous):

lmaooo,

OpenStudy (mathmate):

If you want to know how to do it mentally, it is 28^2-4^2=784-16=768

OpenStudy (anonymous):

yea that works but if you simply make it 25 * 32 which is simply 800 than subtract 32 its a lot easier to do in your head

OpenStudy (mathmate):

That's quite true. I memorize/calculate all the squares in my head between 1 to 110, so that's not much of a problem. They're useful for estimating square-roots also.